OpenStudy (anonymous):

An astronaut weighs 104 N on the moon, where the acceleration of gravity is 1.61 m/s2. How much does he weigh on earth?

OpenStudy (stormfire1):

First, figure out the person's mass: \[\frac{104N}{1.61 m/s^2}= ~64.6kg\]Since 1kg = ~2.2lbs, they will weigh about 142lbs on Earth

OpenStudy (anonymous):

the computer is not accepting my units as lbs or kg,that's my problem what other unit can be used?

OpenStudy (stormfire1):

you can also input it as Newtons: 9.8m/s^2 * 64.6kg = ~633N

OpenStudy (anonymous):

yeah correct it has accepted it,thanks alot

OpenStudy (stormfire1):

np